OpenStudy (taylovesmatthew):

Find the percent of increase from 300 to 390. Round to the nearest tenth of a percent if necessary.

OpenStudy (anonymous):

390/300 = x/100. cross multiply 39000 = 300x divide by 300 and that will give you your answer